OpenStudy (mtalhahassan2):

1/2(x square +1) - 3/2 (x square -1)

OpenStudy (kkutie7):

is this your equation?

OpenStudy (kkutie7):

@MTALHAHASSAN2

OpenStudy (mtalhahassan2):

think so

OpenStudy (mtalhahassan2):

no

OpenStudy (kkutie7):

ok so idk how its written. i need that first

OpenStudy (kkutie7):

\[\frac{1}{2}(x^{2}+1)-\frac{3}{2}(x^{2}-1)\] this then?

OpenStudy (mtalhahassan2):

exactly the same

OpenStudy (mtalhahassan2):

dear

OpenStudy (kkutie7):

*sigh* what is your equation there so i can help?

OpenStudy (mtalhahassan2):

I am confuse what to do with this

OpenStudy (mtalhahassan2):

what happen

OpenStudy (mtalhahassan2):

same equation

OpenStudy (mtalhahassan2):

which one you wroted

OpenStudy (kkutie7):

so the one i type works?

OpenStudy (mtalhahassan2):

yeah

OpenStudy (kkutie7):

\[\frac{1x^{2}}{2}+\frac{1}{2}-\frac{3x^{2}}{2}+\frac{3}{2}\]

OpenStudy (kkutie7):

\[\frac{1x^{2}-3x^{2}+1+3}{2}\]

OpenStudy (mtalhahassan2):

ok

OpenStudy (mtalhahassan2):

i goted

OpenStudy (kkutie7):

\[\frac{-2x^{2}+4}{2}\]

OpenStudy (mtalhahassan2):

thanks dear again

OpenStudy (kkutie7):

\[\frac{-x^{2}+2}{1}\] or \[-x^{2}+2\]